# ALIGNING AUDIO CAPTIONS WITH HUMAN PREFERENCES

### v1

@@ 5 Results and Discussion
We evaluate our RLHF framework on both public and proprietary datasets, including challenging cases where baseline models fail to produce accurate or natural captions. Captions are assessed using n-gram-based metrics, semantic-aware metrics, our custom reward model, and human preference judgments. Table 1 shows that even though the baseline model performs well on automatic metrics, it fails to align with human preferences highlighting the limitations of rule-based and semantic metrics. In contrast, our RLHF model consistently improves human preference win-rates, except on the AudioCaps evaluation set, likely due to limited preference data for training the reward model. Preference based reward model correlates well with human judgments, and RLHF maintains competitive scores on automatic metrics. Despite being built upon the $CLAP$ model, our reward model exhibits a more consistent similarity trend with human evaluations than the conventional $CLAP_{Audio\text{-}Text}$ score. Human preference win-rate analysis reveals that when the baseline model fails to generate high-quality captions, the reward model assigns lower scores to those samples, thereby steering the learning process toward outputs that better reflect human judgment. Table 2 shows two scenarios where proposed RLHF aligns well with humans either based on naturalness or the correctness of the caption, compared to the other methods.
5.1 Ablation study
Table 3 illustrates that increasing the amount of preference data leads to a more effective reward model, which in turn enhances RL fine-tuning. This improvement results in higher-quality caption generation that aligns more closely with human preferences.
Fig. 4: SFT v/s RLHF comparison
Figure 4 compares different training methodologies. For the AudioCaps dataset, supervised fine-tuning was performed using additional ground-truth captions from the train split, while proposed RLHF only uses audio from the same AudioCaps training data. In contrast, for the proprietary dataset, we used the winning captions from the preference dataset as ground-truth for supervised tuning and only audio for the RLHF fine-tuning. Due to limited human annotaotrs, we conducted human study with 2 annotators asked to annotate 100 samples for each experiment. The results indicate that supervised fine-tuning benefits from the availability of paired ground-truth data, leading to improved performance. However, the proposed RLHF approach remains competitive, achieving comparable results despite being fine-tuned without access to additional ground-truth annotations.

### v2

@@ 5 Results and Discussion
We evaluate our RLHF framework on both public and proprietary datasets, including challenging cases where baseline models fail to produce accurate or natural captions. Captions are assessed using n-gram-based metrics, semantic-aware metrics, our custom reward model, and human preference judgments. Table 1 shows that although the baseline model performs well on automatic metrics, it does not align with human preferences, highlighting the limitations of rule‑based and semantic evaluation metrics. In contrast, our RLHF model consistently improves the quality of generated captions, as reflected by higher human preference win‑rates and higher scores from our custom reward model, except on the AudioCaps evaluation set. This exception is likely due to the limited amount of preference data available for training the reward model on that domain. Table 3 illustrates that increasing the amount of preference data leads to a more effective reward model, which in turn enhances RL fine-tuning. This improvement results in higher-quality caption generation that aligns more closely with human preferences. The preference based reward model correlates well with human judgments, and RLHF maintains competitive scores on automatic metrics. Despite being built upon the $CLAP$ model, our reward model exhibits a more consistent similarity trend with human evaluations than the conventional $CLAP_{Audio\text{-}Text}$ score. Human preference win-rate analysis reveals that when the baseline model fails to generate high-quality captions, the reward model assigns lower scores to those samples, thereby steering the learning process toward outputs that better reflect human judgment. Table 2 shows two scenarios where caption generated by RLHF model align well with human written ground-truth captions either based on the naturalness or the correctness of the captions, compared to the other methods.